OpenStudy (anonymous):

Find the value of each expression. 9P3 & 10C4

Parth (parthkohli):

\(n{\rm P}r\) is the number of ways you could take \(r\) objects from \(n\) objects and then arrange them in a line, also known as permutations.

Parth (parthkohli):

\[\dfrac{n!}{(n - r)!} = n{\rm P}r\]Consider the example:\[\rm 9P5 = \dfrac{9!}{(9 - 5)!} = \dfrac{9!}{4!} = 9\cdot 8\cdot 7 \cdot 6\cdot 5\]

OpenStudy (anonymous):

ok thanks

Parth (parthkohli):

You're welcome. :-)